40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 21 12:46:22 2019, gautam, Update, LSC, CM board study YARM_CMslow.pdf
    Reply  Thu Nov 21 13:14:33 2019, Koji, Update, LSC, CM board study 
       Reply  Thu Nov 21 19:08:58 2019, gautam, Update, LSC, High BW lock of Y arm length to PSL frequency highBW_POY.pdfCM_UGF.pdfIN2_ACcoupling.pdf
Message ID: 15044     Entry time: Thu Nov 21 19:08:58 2019     In reply to: 15043
Author: gautam 
Type: Update 
Category: LSC 
Subject: High BW lock of Y arm length to PSL frequency 

Summary:

The Y arm cavity length was locked to the PSL frequency with ~26kHz UGF, and 25 degrees phase margin. Slow actuation was done on ETMY using CM_Slow as an error signal, while fast actuation was done on the IMC error point via the IN2 input of the IMC servo board. Attachment #1 shows the comparison of the in-loop error signal spectra with only slow actuation and with the full CM loop engaged.

Details:

  1. LSC enable OFF.
  2. Configure the CM board for locking:
    • CM board IN1 gain = 25dB.
    • CM_Slow whitening gain = +18dB, make sure the offsets are correctly set. CM_Slow filter bank = -0.015.
    • CM_Slow-->YARM matrix element in LSC input matrix is -2.5.
    • YARM-->ETMY matrix element in LSC Output matrix is 1.
    • AO gain set to +5dB. IMC Servo board IN2 gain starts at -32dB, the path is disabled. The polarity is Plus.
    • Usual YARM FM triggers are set (FM1, FM2, FM3, FM6, FM8), usual YARM servo gain is used (0.01), usual triggering conditions (ON @ TRY>0.3, OFF @ TRY < 0.1), usual power normalization by TRY.
  3. Enable LSC mode, wait for the arm to acquire lock.
  4. Once the digital boosts are engaged, enable the IMC IN2 path, ramp up the gain to -2 dB. Note that this IN2 path is AC coupled, according to this elog. The corner frequency is 1/2/pi/2e3/6.8uF ~11 Hz. This was confirmed by measurement, see Attachment #3. I couldn't find a 2-pin LEMO-->BNC adaptor so I measured at the BNC connector for the IN2 input, which according to the schematic is shorted to the LEMO (which is what we use for the AO path).
  5. Enable the CM board boost.
  6. Ramp up the CM board IN1 gain to +31dB. In this config, the CM_Slow signal is ~18,000 cts pk (with the +18dB whitening gain), so not saturating the ADC.
  7. Ramp up the IMC IN2 gain to 3dB, engage 2 Super Boosts (can't turn on the third). Limiter is always ON.
  8. Use the CM board error point offset adjust to zero the POY11_I error signal average value - there seems to be some offsets when engaging the boosts. The value I used was 0.9 V (this is internally divided by 40 on the CM board).
  9. Whiten the CM_Slow signal - this doesn't seem to have any impact on the noise anywhere.

I hypothesize that the high-frequency noise (>100 Hz) is higher for POY than POX in Attachment #1 because I am using the "MON" port of the demod board - this has a gain of 2, and there could also be some flaky components in this path, hence the high frequency noise is a factor of a few greater in the POY spectrum relative to the POX spectrum (which is using the main demodulated output). For REFL11, we have a low noise preamp generating the input signal so I don't think we need to worry about this too much.

The PC Drive RMS didn't look any stranger than it usually does for the duration of the lock.

Attachment #2 shows the OLTF of the locking servo with the final gains / settings, which are in bold. The loop is maybe a bit marginal, could possibly benefit from backing off one of the super boosts. But the arm has stayed locked for >1 hour.

The purpose of this test was to verify the functionality of the CM board and also the IN2 of the IMC servo board in a low-pressure environment. Once I confirm that the modelled OLTF lines up with the measured, I will call this test a success, and move on to looking at REFL11 in the arms on ALS, PRMI on 3f config. I am returning the REFL11 signal to the input of the CM board, but the SR785 remains hooked up.

Unrelated to this work - PMC alignment was tweaked to improve input power to IMC by ~5%.

Attachment 1: highBW_POY.pdf  59 kB  | Hide | Hide all
highBW_POY.pdf
Attachment 2: CM_UGF.pdf  152 kB  | Hide | Hide all
CM_UGF.pdf
Attachment 3: IN2_ACcoupling.pdf  146 kB  Uploaded Sat Nov 23 12:41:24 2019  | Hide | Hide all
IN2_ACcoupling.pdf
ELOG V3.1.3-